# Search-Based Software Engineering in the Landscape of AI Foundation Models

### v1

@@ Abstract
Search-based software engineering (SBSE), at the intersection of artificial intelligence (AI) and software engineering, has been an active area of research for about 25 years. It has been applied to solve numerous problems across the entire software engineering lifecycle and has demonstrated its versatility in multiple domains. With the recent advancements in AI, particularly the emergence of foundation models (FMs), the evolution of SBSE alongside FMs remains undetermined. In this window of opportunity, we propose a research roadmap that articulates the current landscape of SBSE in relation to foundation models (FMs), highlights open challenges, and outlines potential research directions for advancing SBSE through its interplay with FMs. This roadmap aims to establish a forward-thinking and innovative perspective for the future of SBSE in the era of FMs.
Keywords AI  $\cdot$ Search-based Software Engineering  $\cdot$ Metaheuristic Optimization  $\cdot$ Foundation Models  $\cdot$ Large Language Models  $\cdot$ Vision Language Models  $\cdot$ Multimodal Models
@@ 1 Introduction
Search-based software engineering (SBSE) has been a prominent field for nearly a quarter-century—approaching its silver jubilee—since it was first introduced by Harman and Jones [17] in 2001. Over the years, it has rapidly evolved to address emerging and complex software engineering problems. It has been successfully applied throughout the software engineering (SE) lifecycle, including requirements engineering, software design, software development, software testing, deployment, and maintenance [19].
@@
SBSE uses metaheuristic search optimization techniques to solve SE problems. This involves reformulating SE problems as metaheuristic search problems by creating solution representation, defining fitness functions, and selecting search operators [17]. The problem formulated in this way is then solved by applying search and optimization techniques to identify optimal solutions. These techniques include single-objective, multi-objective, and many-objective search algorithms, each designed for different problem types and levels of complexity [18, 19].
@@
Recently, artificial intelligence (AI) foundation models (FMs) have brought a major transformation, attracting widespread interest across research communities, academia, and industry [23]. These models are trained on vast amounts of data and are capable of performing a wide range of analytical tasks. For specific application domains, these models can be adapted using techniques such as fine-tuning and prompt engineering. Depending on their modalities, architecture, and application domains, FMs can be categorized as follows: Large Language Models (LLMs) like GPT for textual content; Vision Language Models (VLMs) like ResNet for image and video data; Speech and Audio Models (SAMs) like WaveNet for tasks related to speech recognition, synthesis, and audio analysis; and Multimodal Models (MMs) like CLIP that integrate multiple types of data, such as text and images, to enable content generation and comprehension across different modalities [44]. Their ability to generalize knowledge across multiple domains and adapt to diverse applications has enabled new AI-driven innovations, making them a focal point of exploration. As a result, researchers are actively exploring their potential, academic institutions are integrating them into curricula, and industries are leveraging them to enhance automation, decision-making, and user experiences.
Considering rapid AI transformations driven by FMs, in this paper, we present a strategic roadmap exploring the synergistic interplay between SBSE and FMs in the context of future innovations and advancements. We structure this roadmap to explore four key dimensions, as illustrated in Figure 1: (i) how FMs can be leveraged for SBSE design (Section 2), (ii) how FMs can complement SBSE applied for SE problems (Section 3), (iii) how well-studied SBSE practices can be adapted to FMs customized for particular SE activities (Section 4), and (iv) how SBSE and FMs operate together and their synergistic potential (Section 5). For each dimension, we review the current state of the art, discuss open challenges, and present opportunities with a forward-looking vision that outlines potential research directions.
Figure 1: Research roadmap dimensions outlining the potential integration of SBSE with AI foundation models.
2 Foundation Models for Automating Search-based Software Engineering Design
@@
Recently, LLMs have attracted significant interest in evolutionary computation due to their ability to enhance various aspects [42]. An important step in this regard is to use LLMs as search operators for single and multi-objective optimization, such as language model crossover (LMX) for single-objective [28] and LLM-assisted offspring generation for multi-objective optimization [41]. Another line of work is evolutionary algorithm generation and improvement using LLMs. Notable works include using LLMs to generate swarm intelligence algorithms [31], designing heuristic algorithms [24], and evolving search algorithms [45]. In addition, some studies also focused on using LLMs to assist novice users in interpreting and comprehending evolutionary algorithm-generated solutions [25, 38]. While LLM-based evolutionary computation has recently gained attention in the literature, its effectiveness within the SBSE context remains understudied. Moreover, given LLMs’ potential to enhance evolutionary computation, we anticipate that other FMs, such as VLMs and MMs, could further complement SBSE in solving complex SE problems. Considering this, we outline key research opportunities as follows:
2.1 Design of Fitness Functions for SBSE Problems
@@
Typically, a software engineer manually identifies and defines the fitness functions for SBSE problems based on their understanding of the domain. FMs have the potential to assist in automating the definition and subsequent implementation of fitness functions for a given SBSE problem. For example, in test case prioritization, FMs can help provide alternative fitness function options to guide search algorithms in prioritizing test cases, e.g., based on historical data on faults found by each test case and the coverage achieved. We foresee the following options to be investigated with FMs for designing fitness functions for SBSE problems: 1) FMs can be studied as a chatbot created for SBSE users, where they can build fitness functions by interacting with the chatbot and then implement their fitness functions themselves; 2) SBSE users can provide a set of requirements and ask an FM to implement the fitness function. In addition, VLMs, SAMs, and MMs can be used to derive fitness function suggestions from images, audio, and video data, for example, when generating test scenarios for autonomous driving systems (ADS).
2.2 Design of Search Operators
In a recent study, Wu et al. [42] categorized various LLM-based search operators, such as crossover and mutation, for single and multi-objective algorithms. However, the application of such operators in SBSE remains undetermined. Therefore, one potential research direction is to explore LLM-based search operators in the SBSE context. Another research direction is to define search operators using other FMs like VLMs and MMs to enable SBSE applications to emerging domains such as autonomous systems and quantum computing. For example, when dealing with ADS data in image format, VLM-based search operators can be employed to generate test scenarios. Similarly, for ADS data in audio or video formats, search operators based on SAMs and MMs can enhance test scenarios considering the mixed-mode nature of data.
2.3 Solution Encoding for SBSE Problems

### v2

# Search-Based Software Engineering and AI Foundation Models: Current Landscape and Future Roadmap

@@ Abstract
Search-based software engineering (SBSE), which integrates metaheuristic search techniques with software engineering, has been an active area of research for about 25 years. It has been applied to solve numerous problems across the entire software engineering lifecycle and has demonstrated its versatility in multiple domains. With recent advances in AI, particularly the emergence of foundation models (FMs) such as large language models (LLMs), the evolution of SBSE alongside these models remains undetermined. In this window of opportunity, we present a research roadmap that articulates the current landscape of SBSE in relation to FMs, identifies open challenges, and outlines potential research directions to advance SBSE through its integration and interplay with FMs. Specifically, we analyze five core aspects: leveraging FMs for SBSE design, applying FMs to complement SBSE in SE problems, employing SBSE to address FM challenges, adapting SBSE practices for FMs tailored to SE activities, and exploring the synergistic potential between SBSE and FMs. Furthermore, we present a forward-thinking perspective that envisions the future of SBSE in the era of FMs, highlighting promising research opportunities to address challenges in emerging domains.
Keywords AI $\cdot$ Search-based Software Engineering $\cdot$ Metaheuristic Optimization $\cdot$ Evolutionary Computation $\cdot$ Foundation Models $\cdot$ Large Language Models $\cdot$ Vision Language Models $\cdot$ Multimodal Models
@@ 1 Introduction
Search-Based Software Engineering (SBSE), formally introduced as a field by Harman and Jones [66] in 2001, has gained prominence over the past quarter century, although the use of search algorithms to address software engineering (SE) problems dates back to earlier work, such as test data generation by Miller and Spooner [115] in 1976. Over the years, it has rapidly evolved to address emerging and complex SE problems. It has been successfully applied throughout the SE lifecycle, including requirements engineering, software design, software development, software testing, deployment, and maintenance [68].
@@
SBSE applies artificial intelligence (AI) techniques, including evolutionary computation and metaheuristic search, to effectively solve SE problems. This involves reformulating SE problems as metaheuristic search problems by creating solution representation, defining fitness functions, and selecting search operators [66]. The problem formulated in this way is then solved by applying search and optimization techniques to identify optimal solutions. These techniques include single-objective, multi-objective, and many-objective search algorithms, each designed for different problem types and levels of complexity [67, 68].
@@
Recently, foundation models (FMs)—a term introduced by Bommasani et al. [19]—have brought a major transformation, attracting widespread interest across research communities, academia, and industry [91, 206]. Although the term “foundation models” broadly refers to large-scale pre-trained models, it is most commonly associated with large transformer-based architectures, such as GPT, BERT, and CLIP [19]. These models are trained on vast amounts of data and are capable of performing a wide range of analytical tasks. For specific application domains, these models can be adapted using techniques such as fine-tuning and prompt engineering. Depending on their modalities, architecture, and application domains, FMs can be categorized as follows: Large Language Models (LLMs) like GPT for textual content; Vision Language Models (VLMs) like ResNet for image and video data; Speech and Audio Models (SAMs) like WaveNet for tasks related to speech recognition, synthesis, and audio analysis; and Multimodal Models (MMs) like CLIP that integrate multiple types of data, such as text and images, to enable content generation and comprehension across different modalities [201]. Their ability to generalize knowledge across multiple domains and adapt to diverse applications has enabled new AI-driven innovations, making them a focal point of exploration. As a result, researchers are actively exploring their potential, academic institutions are integrating them into curricula, and industries are leveraging them to enhance automation, decision-making, and user experiences [168].
Considering rapid AI transformations driven by FMs, in this paper, we present a research roadmap exploring the synergistic interplay between SBSE and FMs. We first analyze the current research landscape to examine the directions in which SBSE and FMs are being integrated and the types of SE problems that are being addressed. From this analysis, we identify open challenges and opportunities for developing novel techniques that combine SBSE and FMs to solve complex domain-specific and SE problems. We also highlight the challenges of conducting empirical evaluations of such techniques, which require significant attention from the research community. Finally, we outline a 2030 research horizon, using McLuhan’s Tetrad [111]—a framework designed to analyze the effects of new technologies—to illustrate how FMs, as disruptive technology, can shape the future of SBSE. Furthermore, we present research opportunities for potential synergy between SBSE and FMs to address problems in emerging domains.
2 Roadmap Overview
Figure 1 illustrates the overall structure of the roadmap.
Figure 1: Roadmap overview showing the discussion flow: strengths and weaknesses (Section 3), SBSE–FM synergies (Sections 4, 5, 7, 6 and 8), empirical evaluation challenges (Section 9), and the 2030 research horizon (Section 10).
We begin by describing the strengths and weaknesses of the SBSE and the FMs (Section 3). Following this, we explore the potential synergy between SBSE and FMs (as shown in Figure 2), analyzing how the strengths of one can complement the other to advance SE (Sections 4, 5, 7, 6 and 8). Subsequently, we discuss the challenges associated with conducting empirical evaluations using techniques that integrate SBSE and FMs (Section 9). Finally, we outline the research horizon for 2030, including open research challenges and opportunities ahead (Section 10).
Figure 2 shows five aspects of the potential synergy between SBSE and FMs.
Figure 2: Key aspects of the potential synergy between SBSE and FMs. The abbreviations used are: FMs (Foundation Models), LLMs (Large Language Models), VLMs (Vision-Language Models), $M$s (Multimodal Models), SBSE (Search-Based Software Engineering), SE (Software Engineering), SDLC (Software Development Lifecycle), and ADS (Autonomous Driving Systems).
Through a review of the recent literature, we identified five core aspects that characterize how the research community integrates FMs with SBSE. These aspects focus on research questions: (1) how FMs can be leveraged for SBSE design (Section 4), (2) how FMs can complement SBSE applied to SE problems (Section 5), (3) how SBSE can be applied to deal with the challenges of FMs (Section 6), (4) how well-studied SBSE practices can be adapted to FMs customized for particular SE activities (Section 7), and (5) how SBSE and FMs operate together and their synergistic potential (Section 8). For each aspect, we provide a review of current state-of-the-art research, followed by a roadmap with challenges and research opportunities.
3 Strengths and Weaknesses of SBSE and FMs
In this section, we begin by exploring the advantages and limitations of SBSE, followed by an analysis of the benefits and downsides of FMs.
3.1 Advantages and Limitations of SBSE
In SBSE [66], SE problems are cast into either search or optimization problems. Then, any kind of search or optimization algorithm can be applied to attempt to solve those SE problems.
On the one hand, given $S$, the space of all possible solutions for a problem $X$, a “search” problem is when we need to find one solution $s\in S$ for which a boolean predicate is evaluated as true, i.e., $f(s)=true$. For example, $X$ could be the SE problem of automated bug fixing [6, 72], where $S$ is the space of all possible patches, and $f(s)$ verifies if, for the given patch $s$, all current test cases pass (and so the bug can be considered as fixed).
On the other hand, in an “optimization” problem, the function $f(s)$ gives a numerical score. Depending on the problem, the objective would be to either minimize or maximize such a score. For example, in test data generation [112, 2], where $S$ is the space of all possible test cases for the given software under test, then $f(s)$ could be a metric like code coverage, where we try to find the test case $s_{m}\in S$ with maximum value for $f$.
For any non-trivial problem, the space $S$ of all possible solutions is extremely large. Given $n$, the length in bits of the representation of a solution, there are $2^{n}$ possible combinations. This exponential number of combinations cannot be brute forced, i.e., enumerated fully. For example, consider something as simple as the testing of the function classify(a,b,c) to classify triangles from the length of their edges (a common example in the software testing literature [120, 112, 11]). There are four possible outcomes: scalene, isosceles, equilateral, and invalid. Given the three inputs as 32-bit integers, there are $|S|=2^{96}$ possible inputs for this function. That is more than 79 octillion possible combinations.
Given a reasonable amount of time, only an extremely tiny subset of $S$ can be evaluated using the function $f$. Brute force approaches in which each solution in $S$ is enumerated and evaluated with $f$ are infeasible in all but trivial problems. A search algorithm will evaluate some solutions in $S$, compute their score $f(s)$, and then decide which other solutions in $S$ to evaluate. Such a decision is made with the so-called search operators (e.g., do small random modifications to an already evaluated solution with a good $f$ score). This process is then repeated in a loop, until either $k$ solutions have been evaluated, or a certain amount of time (e.g., 1 hour) has passed (i.e., based on the chosen stopping criterion). The longer a search algorithm is left exploring the search space $S$ by evaluating solutions in it using $f$, the better results can be expected.
There are many different search algorithms, including, for example, Hill Climbing [158], Genetic Algorithms [71], and Ant-Colony Optimization Algorithms [41]. However, no algorithm is best on all possible problems [190]. Mathematically, it is not possible. The more domain knowledge specific to a problem $X$ can be exploited by a search algorithm, the better results can be expected. This is one of the main drivers for designing ad-hoc, specific search algorithms tailored for the specific properties of the addressed problem $X$. An example is the Many Independent Optimization Algorithm for test suite generation [7].
There are many problems in science and engineering that can be modeled as search or optimization problems [125, 86]. Traditional problems include, for example, the Traveling Salesperson Problem [142] and the Knapsack Problem [143]. But they appear in all domains, from bioengineering to economics. SE is just one of many kinds of fields that has tasks that can be modeled as search/optimization problems. There are decades of success stories about the application of search/optimization algorithms [125, 86]. Their success in addressing SE problems is, therefore, not unexpected [68].
As for any technique, there are places and contexts in which search algorithms do not work well. One of the major challenges in applying search to SE problems is the definition and design of the score function $f$. This is problem-dependent and requires expertise to create one. For example, assume the problem $X$ of generating test cases. On the one hand, you can try to optimize the generation of test cases to maximize metrics such as code coverage. This can be implemented in a $f$ function relatively easily (most programming languages already have existing libraries to measure code coverage). But what about code readability? I.e., what about the generation of test cases that are easy to understand for humans? How to mathematically quantify in a function $f$ if a piece of code $s$ is easy to read and understand? If you cannot design and implement a score function $f$ for a given problem $X$, then you cannot use SBSE to address such a problem. For several tasks in SE, where, for example, natural language is involved (e.g., comments in Pull Requests, or text descriptions of product requirements), it might be challenging, if possible at all, to define a score function $f$.
Another major issue is that, to be a viable option, the score function $f$ must not be too slow and resource-intensive to compute. A search algorithm, to provide good enough results, might need to evaluate tens or hundreds of thousands of solutions $s$. This, of course, depends on how long the search process is left running (i.e., depending on the chosen stopping criterion). For example, in the literature, experiments on unit test generation are run for something like 2 minutes [52], whereas experiments on system testing are run for 1 to 24 hours [59]. If the score function $s$ only needs a few milliseconds to compute, then in these contexts the use of SBSE is feasible (as many success stories have shown [68]). However, what about the cases in which $f$ might take hours to compute? Even if $f$ takes “only” 1 minute to compute, at most only 1440 solutions could be evaluated in a single day. Furthermore, what about all other cases in which a response (i.e., finding best $s$) must be provided quickly, e.g., in the order of a few (milli)seconds? In all those cases, likely SBSE techniques would not be particularly beneficial.
When talking about the execution time of $f$, there are two important aspects to consider: parallelization and real-world constraints. Advanced search algorithms are able to compute the score function $f$ in parallel. Instead of evaluating one solution $s$ at a time, a search algorithm could evaluate $n$ in parallel, where $n$ could be based on the number of CPUs available on the machine the algorithm is running. Of course, this will impact how search operators are defined, but there are several different techniques to handle it (e.g., “island models” [189]). Based on the hardware available, parallelization can provide a linear speed-up to the search process. The possibility to automatically adapt to the available hardware resources is a major feature of search algorithms.
But, how much hardware is available? This depends on the real-world scenarios in which these techniques are applied. In the case of SE problems, a common example is developers and testers using their work laptops for SBSE tasks. As of 2025, a typical laptop can have several CPU cores (e.g., not uncommon to have between 4 and 8 CPU cores). For example, this very text is written on a Mac with an M4 chip having a 16-core CPU. Considering the cost of hardware (e.g., laptops and servers) and electricity, compared to the salary cost of professional software engineers, the use of extra hardware dedicated to SBSE tasks can be warranted. This is the case if the use of SBSE brings a positive return on investment by saving (the expensive) time of software engineers.
Without disrupting existing software development processes, extra hardware that could be used for SBSE tasks might already be available. For example, considering a common practice like the use of Continuous Integration (CI) servers [159], extending such servers to run SBSE tasks is possible (e.g., as done in [27, 12]). Different SE tasks, like test generation, bug fixing, and so on, could automatically be run in CI after developers make code changes.
In summary, SBSE can excel in all problems where a score function $f$ can be defined, and its computational cost is not too high. The more a search process is left running, the better results can be obtained. SBSE techniques can be successfully run on developer laptops, as well as being able to exploit high parallelism when run on dedicated servers. Their adaptability to available hardware is one of their major strengths.
3.2 Benefits and Downsides of FMs
FMs represent a groundbreaking advancement in AI due to their ability to process multimodal data [19]. These models encompass diverse architectures, such as LLMs for natural language processing and VLMs or MMs for handling image, audio, and video data [201]. A key strength of FMs is that they are trained using enormous and heterogeneous datasets, due to which they possess extensive cross-domain knowledge and the ability to identify complex patterns across different types of inputs. FMs also feature task versatility due to their transformer-based architectures and self-supervised learning techniques. These capabilities enable them to understand context and provide outputs with reasoning that closely resemble human thinking, making them highly effective for complex decision-making and problem-solving tasks.
One of the most notable benefits of FMs is their adaptability and generalization in various domains. Unlike traditional models that require extensive retraining for new tasks, FMs can adapt and generalize to domain-specific tasks with minimal fine-tuning. This adaptability has driven their adoption in numerous fields, including SE, autonomous systems, and robotics [70, 48, 57]. A recent advancement is vision–language–action (VLA) models [145], which extend the multimodal capabilities of FMs by not only interpreting complex inputs, but also generating executable actions. This makes VLAs particularly valuable in dynamic and interactive environments, such as autonomous vehicles or robotics, where they enable advanced control and embodied intelligence [145].
Despite their significant benefits, FMs also come with certain downsides. A primary concern is that these models often struggle in handling dynamic systems, as they lack the real-time adaptability needed to respond to evolving requirements or changes in feedback loops [48]. For example, in applications such as autonomous systems, FMs lack the ability to dynamically incorporate new information. Furthermore, their inherent limitations, such as hallucination, non-determinism, uncertainty, and output variability, raise serious reliability concerns, especially in safety-critical domains such as robotics and healthcare, where factually incorrect outputs may cause safety risks [29, 80].
Another significant challenge lies in the resource-intensive nature of FMs’ fine-tuning and retraining for specific applications. Customizing these models requires substantial computational resources, access to high-quality domain-specific data, and expertise, which may not be feasible for all applications [119]. Moreover, the rapid evolution of FMs requires frequent updates and integrations for applications built on them, which pose challenges to long-term stability and compatibility [138]. In industrial settings, where trust and security are key concerns, FMs can introduce risks related to sensitive data and schema-related vulnerabilities, which can limit their adoption [138]. Finally, ethical considerations of FMs, such as bias, fairness, and the potential for misuse, further complicate their deployment in socially sensitive applications like finance [119]. These challenges emphasize the need for ongoing research to address the limitations of FMs and ensure their safe and responsible deployment.
4 FMs for SBSE Design
In this section, we discuss the potential of FMs to advance SBSE design, starting with an overview of the current state of research, followed by a discussion of the key challenges and opportunities for further research.
4.1 Current Landscape
The typical SBSE process comprises several key steps, including problem formulation, solution encoding, search operator selection, defining the search space, designing the fitness function, selecting the search algorithm, and specifying the stopping criteria [66]. Each of these steps requires careful consideration when applied to SE problems. Over the years, a substantial body of research has focused on the design of SBSE solutions for various phases of the SE lifecycle, from requirements engineering to testing and deployment [33, 10]. For example, in requirements engineering, several efforts have been devoted to developing appropriate solution encoding schemes, designing specialized search operators, and crafting fitness functions to optimize requirements engineering problems such as requirement selection, prioritization, fairness analysis, and traceability [68]. For software design, the available literature has focused on various aspects such as defining search operators and fitness functions for model and architecture optimization [24, 28]. Another area that has gained significant attention is search-based software testing [113]. In this area, research has focused on devising novel mutation operators for cost-effective multi-objective test case selection [176], automatically inferring fitness functions for test case generation [49], dynamically selecting fitness functions for unit test generation with high fault detection effectiveness [4], reducing search space and reusing previous solutions for test data generation [148, 155], and improving quality indicators for multi-objective optimizations such as test selection [157, 130]. Similarly, in software debugging, several studies have focused on various aspects, including the selection of mutation operators and the refinement of fitness functions to enhance the effectiveness of automated program repair [63, 65, 72, 123].
Recently, LLMs have attracted significant interest in evolutionary computation due to their ability to enhance various aspects [193]. An important step in this regard is to use LLMs as search operators for single and multi-objective optimization, such as language model crossover (LMX) for single-objective [114] and LLM-assisted offspring generation for multi-objective optimization [187]. In the same line of work, Freire et al. [54] investigated the use of LLMs to generate customized suggestions for objective functions aimed at optimizing product line architecture designs.
@@ 4.1 Current Landscape
Another line of work is evolutionary algorithm generation and improvement using LLMs. Notable works include using LLMs to generate swarm intelligence algorithms [132], designing heuristic algorithms [98], and evolving search algorithms [202]. In addition, some studies also focused on using LLMs to assist novice users in interpreting and comprehending evolutionary algorithm-generated solutions [108, 161]. While LLM-based evolutionary computation has recently gained attention in the literature, its effectiveness within the SBSE context remains understudied. Given LLMs’ potential to enhance evolutionary computation, we anticipate that other FMs, such as VLMs and MMs, could further complement SBSE in solving complex domain problems like autonomous systems [163, 57], robotics, and quantum computing [118]; for example, in the autonomous driving domain, LLMs have been combined with search algorithms to generate tests: LLMs first derive logical scenarios from reports of real accidents, and then search algorithms search for concrete scenarios in the search space defined by these logical scenarios [171].
4.2 Roadmap
Although the recent literature predominantly focuses on using LLMs (rather than the broader scope of FMs) for various aspects of SBSE design, numerous challenges persist, which present valuable opportunities for further research. In the following, we explore these open challenges and research opportunities in detail.
4.2.1 Fitness Functions for SBSE Problems
@@
Typically, a software engineer manually identifies and defines the fitness functions for SBSE problems based on their understanding of the domain. FMs have the potential to assist in automating the definition and subsequent implementation of fitness functions for a given SBSE problem. For example, in test case prioritization, FMs can help provide alternative fitness function options to guide search algorithms in prioritizing test cases, e.g., based on historical data on what type of tests are more likely to detect faults. We foresee the following options to be investigated with FMs for designing fitness functions for SBSE problems: (1) FMs can be adapted as chatbots for SBSE users, who can build fitness functions by interacting with the chatbots and then implement the fitness functions themselves; (2) SBSE users can provide a set of requirements and ask the FM to implement the fitness function. In addition, VLMs, SAMs, and MMs can be used to derive fitness function suggestions from images, audio, and video data, for example, when generating test scenarios for autonomous driving systems (ADS).
4.2.2 Solution Encoding for SBSE Problems